The computer that controls a bank's automatic teller machine crashes a mean of 0.4 times per day. What is the probability that, in any seven-day week, the computer will crash at least 4 times? Round your answer to four decimal places.
Answer #1

= 0.4 * 7 = 2.8

It is a Poisson distribution.

P(X > 4) = 1 - P(X < 4)

= 1 - (P(X = 0) + P(X = 1) + P(X = 2) + P(X = 3))

= 1 - (e^(-2.8) * (2.8)^0/0! + e^(-2.8) * (2.8)^1/1! + e^(-2.8) * (2.8)^2/2! + e^(-2.8) * (2.8)^3/3!)

= 1 - 0.6919

= 0.3081